OpenStudy (anonymous):

solve the equation by finding square roots. 3x^=15

OpenStudy (lgbasallote):

that's 3x^ what?

OpenStudy (anonymous):

It is just 3x^=15 and i have to find the answer in square root.

OpenStudy (lgbasallote):

^ means exponent...what's the exponent?

OpenStudy (anonymous):

These are Quadratic Equations. and the question asks us to solve the equation by finding the square roots. 3x squared = 15. that is all i understand.

OpenStudy (anonymous):

x^2=15/3 x^2=5 x= \[\pm \sqrt{5}\]